a boxing ring is in the shape of a square,20 ft on each side. how far apart are the fighters if they are in opposite corners of the ring?
Answer by stanbon(75887)

a boxing ring is in the shape of a square,20 ft on each side. how far apart are the fighters if they are in opposite corners of the ring?
-------------
Draw the picture of the square.
Each side is 20 ft.
Use Pythagoras to solve for the length of the diagonal:
x^2 = 20^2 + 20^2
x^2 = 2*20^2
x = 20*sqrt(2)
x = 28.284... ft
=====================
Cheers,
Stan H.